# Do We Need Subsidiarity in Software?

### v1

@@ 4.2.2. Interview Protocol
The interviews were conducted in person or on Zoom (with cameras turned on) by the first author and took a semi-structured format. Participants were first introduced to the benefits and tradeoffs of data privacy and sharing data with tech companies. Then, for each of the six software platform types — operating systems, browsers, email applications, search engines, social media, and chat applications — participants were led through a thought exercise where they envisioned why it might be necessary for data on those types of applications to be controlled at the platform, community, or user level, as defined in Section 3. The purpose of the thought exercise was to prepare the participants to identify what they think is the necessary level of control for data on that type of software application, and to understand participant motivations for choosing that level of control. The interviewer let the participants come up with their own examples first, but if the participants struggled to come up with examples the interviewer provided some of their own examples in order to help the participant make an informed decision. After the thought exercise, the participant was invited to identify which level of control is the lowest level possible and highest level necessary for that type of software application. Participants distinguished between the cases of personal use and group use (encompassing usage in circumstances such as workplaces, schools, and special interest communities) in their responses.
@@ 4.2.3. Data Analysis
Audio from the interviews was recorded and transcribed. We counted the number of participant responses for desired platform, community, and user levels of control for each software type in both the personal usage and group usage contexts. We then analyzed the interview transcripts using a reflexive thematic analysis approach (Braun and Clarke, 2019) to better understand participant motivations for selecting each level of control. We inductively coded for the reasons participants selected the levels of control that they did for each different software type. We also inductively coded for any nuances participants shared in selecting a level of control. Within each of the four categories — reasons for selecting user level of control, reasons for selecting community level of control, reasons for selecting platform level of control, and nuances in selecting a particular level — we synthesized the codes for that category into overarching themes.
@@ 5. Data Flow Monitoring Results
We note that although we excluded IP addresses that were confirmed background IPs from a particular operating system or browser, in many cases, previously unseen IP addresses from the same organizations as the background traffic popped up only in the experiments on that particular OS or browser. We report these suspected background interactions in the figures and mark them as suspected background interactions. However, we omit these suspected background contacts from the written analysis for clarity and brevity. We also omit the number of times each organization was contacted and total bytes of application data transferred in the written analysis, but include these in the summarizing figures.
@@ 5.1. Idle Operating Systems
In a one hour window, Windows contacted Microsoft, Akamai, Google, our university, and Cloudflare. Of these, application data over 1000 bytes was transferred to Microsoft, Akamai, and Google. Ubuntu contacted Canonical Group (the software company that maintains Ubuntu) and our university, and transferred no application data over 1000 bytes to either. Figure 2 summarizes our idle OS results.
@@ 5.2. Idle Browsers
In a one hour window, Chrome contacted Google, Qwilted (a Content Delivery Network), and Fastly (a Content Delivery Network) on Windows, and Google, Amazon, and Fastly on Ubuntu. For both, Google was the only platform that application data over 1000 bytes was transferred to.
Firefox contacted Google Cloud, Google, and Akamai on both operating systems, and additionally contacted Fastly and Qwilted on Windows. Application data over 1000 bytes was transferred to Google Cloud and Akamai on Windows. In the Akamai case, this application data transfer could have been background from the Windows operating system rather than from Firefox itself due to observed contact with Akamai from the Windows OS.
Brave contacted Amazon on both operating systems, and transferred over 1000 bytes of application data in both cases.
Figure 3 summarizes our idle browser results.
@@ 5.3.3. Social Media:
Instagram contacted Facebook, Three Ireland Limited, and Turk Telecom Times. Of these, Instagram transferred application data to Facebook. TikTok contacted Akamai: although Akamai is part of Firefox background traffic, the number of contacts to Akamai compared to the other idle experiments, as well as the volume of application data transferred to Akamai, indicate that contact to Akamai is from TikTok as well. TikTok only transferred application data to Akamai. X contacted Cloudflare and transferred application data. Bluesky contacted OVH SAS, Amazon, Datacamp-Limited, and Amazon AES (Advanced Encryption Standard). Bluesky transferred application data to OVH SAS. YouTube contacted Google and Indiana GigaPop and did not transfer any application data while running idle. Figure 6 summarizes our social media results.
@@ 5.3.4. Chat:
Slack contacted Amazon, Amazon AES, Rubicon Project, and Cloudflare. Slack transferred application data to Amazon. Discord contacted Cloudflare and Amazon, and did not transfer any application data. Figure 7 summarizes our chat application results.
@@ 5.4.1. Email:
While sending an email, Gmail contacted Google on all browsers, Amazon on Chrome, Google Cloud on Chrome, and Fastly on Brave. Of these, only Amazon on Chrome is unexpected. The University Gmail contacted Google and Amazon on all browsers. All are expected: Google is the parent company, and Amazon is expected because our university Gmail account uses single sign-on with Okta, which runs on AWS. Proton Mail contacted Proton on all browsers and Amazon on Firefox. The contact with Amazon on Firefox would be unexpected, however, Proton Mail uses AWS for alternative routing (pro, [n. d.]), thus, Proton Mail has no unexpected contact. Figure 8 summarizes our results.
[…]
@@ 5.4.2. Search:
Google search contacted Google on all three browsers. This is expected: Google is the parent company. Duck Duck Go contacted Microsoft on all browsers. This is expected: Duck Duck Go is powered by Microsoft Bing. Brave contacted Amazon on all three browsers. This is expected: Brave runs on AWS. Figure 9 summarizes our results.
Figure 9. The organizations contacted while performing a search on Google Search, Duck Duck Go, and Brave Search. No contacts are unexpected.
[…]
@@ 5.4.3. Social Media:
Instagram contacted Facebook on all browsers; this is expected as the parent company. Instagram also contacted Internet service providers in other countries: Getlinks SMC Private Limited (Pakistan), Horizon Scope Mobile Telecom (Iraq), Reliance Jio Infocomm Limited (India), COSCOM Liability Limited Company (Uzbekistan), and Telecom Algeria (Algeria). This is unexpected contact given our location in the Midwestern United States. TikTok contacted Akamai, Google, and Fastly on all browsers. Akamai and Fastly are expected due to content delivery; Google is unexpected. X contacted Cloudflare, Google, and Fastly on all browsers. Cloudflare and Fastly are likely content delivery; Google is unexpected. Bluesky contacted Bluesky, OVH SAS, Datacamp Limited, and Amazon on all browsers. This is expected traffic as the parent company and content delivery network. (Note: Bluesky used to run on Amazon Web Services; now it does not. However given that there could be some lingering contact with AWS due to platform migration, we give Bluesky the benefit of the doubt that its contact with Amazon is not due to unexpected data sharing.) Bluesky additionally contacted Amazon-AES on Brave which is likely for content delivery. YouTube contacted Google and Indiana Gigapop on all browsers. Google is the parent company of YouTube and Indiana Gigapop is a high-speed routing network for academic institutions in the Midwestern United States, so both are expected. Figure 10 summarizes our results.
@@ 5.5. Determining Levels of Control
Figure 12. Determining an operating system’s level of control based on if it sends application data during idle time, communicates with unexpected IPs, and if its parent company is a known data harvester. If any of these three criteria are met, the operating system operates at a higher level of control; if none are met, it operates at a lower level. Windows meets all three criteria thus it operates at a higher level. Ubuntu meets none of the criteria so it operates at a lower level.
[…]
@@
Figure 13. Determining a browser’s level of control based on if it sends application data during idle time, communicates with unexpected IPs, and if its parent company is a known data harvester. If any of these three criteria are met, the operating system operates at a higher level of control; if none are met, it operates at a lower level. Each of the browsers meets at least one of the criteria so they all operate at a higher level of control.
[…]
@@
Figure 14. Determining an application’s level of control based on if it sends application data during idle time, communicates with unexpected IPs, and if its parent company is a known data harvester. If any of these three criteria are met, the operating system operates at a higher level of control; if none are met, it operates at a lower level. Proton Mail, Duck Duck Go, and Brave Search are the only applications that operate at a lower level of control.
[…]
@@ 5.5.2. Browsers:
Chrome contacted Amazon which was unexpected, and transferred application data while running idle. Additionally, Chrome is owned by Google: a known data harvester (Schelter and Kunegis, 2016; Gill et al., 2013; Roesner et al., 2012; Lerner et al., 2016; Karaj et al., 2018; Englehardt and Narayanan, 2016; Krishnamurthy and Wills, 2009; Cassel et al., 2022). Thus, Chrome operates at a higher level of control.
@@ 5.5.6. Chat:
Slack transferred application data while running idle and contacted unexpected IPs, mostly advertising providers, both while running idle and while taking an action. Discord did not transfer application data while running idle however it did unexpectedly contact Microsoft and Google during use. Thus, both applications operate at a higher level of control.
@@ 6.1. What levels of control did participants identify as the necessary level?
6.1.1. Operating Systems:
For personal usage, 12 participants selected the user level, 2 participants selected the community level, and 2 participants selected the platform level. For group usage, 0 participants selected the user level, 15 participants selected the community level, and 1 participant selected the platform level.
6.1.2. Browsers:
For personal usage, 10 participants selected the user level, 2 participants selected the community level, and 4 participants selected the platform level. For group usage, 1 participant selected the user level, 12 participants selected the community level, and 3 participants selected the platform level.
6.1.3. Email:
For personal usage, 13 participants selected the user level, 1 participant selected the community level, and 2 participants selected the platform level. For group usage, 1 participant selected the user level, 13 users selected the community level, and 2 participants selected the platform level.
6.1.4. Search engines:
For personal usage, 8 participants selected the user level, 2 participants selected the community level, and 6 participants selected the platform level. For group usage, 0 participants selected the user level, 11 participants selected the community level, and 5 participants selected the platform level.
6.1.5. Social media:
For personal usage, 11 participants selected the user level, 0 participants selected the community level, and 5 participants selected the platform level. For group usage, 3 participants selected the user level, 10 participants selected the community level, and 3 participants selected the platform level.
6.1.6. Chat:
For personal usage, all 16 participants selected the user level. For group usage, 2 people selected the user level, 13 people selected the community level, and 1 person selected the platform level.
@@ 6.1.6. Chat:
Figure 15. Heat map of user responses for if they wanted user, community, or platform control of their data in different types of platforms for personal usage. The majority of participants wanted user control for all software types. However, more participants wanted platform control for browsers, search engines, and social media.
[…]
@@
A heatmap titled ”What level of control did our participants want in personal usage?” The table has four columns: “User,” “Community,” and “Platform” under the main heading “Level of Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” and “Chat.” A legend on the right indicates the color shading based on the number of participants: Dark Red: 12-16 participants, Red: 8-11 participants, Light Red: 4-7 participants, Pink: 1-3 participants, Gray: 0 participants. Here is a breakdown of the data in the heatmap: OS: “User” column is dark red (12-16 participants). “Community” column is pink (1-3 participants). ”Platform” column is pink (1-3 participants). Browser: “User” column is red (8-11 participants). “Community” column is pink (1-3 participants). “Platform” column is light red (4-7 participants). Email: “User” column is dark red (12-16 participants). “Community” column is pink (1-3 participants). “Platform” column is pink (1-3 participants). Search: “User” column is red (8-11 participants). “Community” column is pink (1-3 participants). “Platform” column is light red (4-7 participants). Social Media: “User” column is red (8-11 participants). “Community” column is gray (0 participants). “Platform” column is light red (4-7 participants). Chat: “User” column is dark red (12-16 participants). “Community” column is gray (0 participants). “Corporate” column is gray (0 participants).
@@
Figure 16. Heat map of user responses for if they wanted user, community, or platform control of their data in different types of platforms for group usage. The majority of participants wanted community control for all software types. However, more participants wanted platform control for search engines.
[…]
@@
A heatmap titled “What level of control did our participants want in group usage?” The table has four columns: “User,” “Community,” and “Platform” under the main heading “Level of Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” and “Chat.” A legend on the right indicates the color shading based on the number of participants: Dark Red: 12-16 participants, Red: 8-11 participants, Light Red: 4-7 participants, Pink: 1-3 participants, Gray: 0 participants. Here is a breakdown of the data in the heatmap: OS: “User” column is gray (0 participants).“Community” column is dark red (12-16 participants). “Platform” column is pink (1-3 participants). Browser: “User” column is pink (1-3 participants). “Community” column is dark red (12-16 participants). “Platform” column is pink (1-3 participants). Email: “User” column is pink (1-3 participants). “Community” column is dark red (12-16 participants). “Platform” column is pink (1-3 participants). Search: “User” column is gray (0 participants). “Community” column is red (8-11 participants). “Platform” column is light red (4-7 participants). Social Media: “User” column is pink (1-3 participants). “Community” column is red (8-11 participants). “Platform” column is pink (1-3 participants). Chat: “User” column is pink (1-3 participants). “Community” column is dark red (12-16 participants). “Platform” column is pink (1-3 participants).

### v2

@@ 4.2.2. Interview Protocol
The interviews were conducted in person or on Zoom (with cameras turned on) by the first author and took a semi-structured format. Rather than asking participants their familiarity with data privacy, which could elicit inconsistent responses on the part of the participants, we instead focused on providing participants with the information about data privacy they needed to make informed responses in the interviews. Participants were first introduced to both the benefits and tradeoffs of data privacy and sharing data with tech companies. The interviewer then shared with the participant the results of our data flow monitoring study to illustrate how common software applications are collecting data, especially in ways that may have been unexpected to the participant. Then, for each of the six software platform types — operating systems, browsers, email applications, search engines, social media, and chat applications — participants were led through a thought exercise where they envisioned why it might be necessary for data on those types of applications to be controlled at the platform, community, or user level, as defined in Section 3. An example of the thought exercise for search engines is as follows. For platform-level control the interviewer asked, Pretend that you’re a software engineer at Google working on the search platform. Why might you think that it would be necessary for Google to have control over user data? For community-level control the interviewer asked, Pretend that you are the head of a company, university or special interest group that has provided laptops to all its members. For example, pretend you are the head of IT at our university or the principal of a high school that has provided laptops to all its students. Why might it be necessary for you, the head of the organization, to control user data from a search engine? For user-level control the interviewer asked Why might it be necessary for the user to have full control over their search data? A full interview guide is included in the appendix.
The purpose of the thought exercise was to prepare the participants to identify what they think is the necessary level of control for data on that type of software application, and to understand participant motivations for choosing that level of control. The interviewer let the participants come up with their own examples first. The interviewer had her own examples on hand if the participant struggled in order to help the participant make an informed decision. For consistency across participants, even if the participant did not struggle to come up with examples, the interviewer shared all her examples after the participant finished brainstorming. Examples of these for search engines included personalization for a better user experience and advertising to make the product free at the platform level of control, blocking unsafe sites from search results at the community level, and privacy, especially privacy for sensitive data, at the user level. Examples for all platforms are included in the aforementioned interview guide. After the thought exercise, the participant was invited to identify which level of control is simultaneously the lowest level possible and highest level necessary for that type of software application. Participants distinguished between the cases of personal use and group use (encompassing usage in circumstances such as workplaces, schools, and special interest communities) in their responses.
@@ 4.2.3. Data Analysis
Audio from the interviews was recorded and transcribed. We counted the number of participant responses for platform, community, and user levels of control for each software type in both the personal usage and group usage contexts. We then analyzed the interview transcripts using a reflexive thematic analysis approach (Braun and Clarke, 2019) to better understand participant motivations for selecting each level of control. The first author inductively coded for the reasons participants selected the levels of control that they did for each different software type. The first author also inductively coded for any nuances participants shared in selecting a level of control. Within each of the four categories — reasons for selecting user level of control, reasons for selecting community level of control, reasons for selecting platform level of control, and nuances in selecting a particular level — the research team synthesized the codes for that category into overarching themes. Specifically, the first author and second author discussed together the codes the first author identified, discussed and debated how to group the codes into themes, and how to most appropriately name the themes to encapsulate the range of codes they encompass.
Positionality Statement: The research team is entirely Christian, white, and American. The first and second authors who performed the reflexive thematic analysis are female. The advising author, who did not participate in the reflexive thematic analysis process, is male. The first author has a bachelors and masters degree in computer science and is pursuing a PhD in computer science. The second author has a bachelors, masters, and PhD in bioengineering, did postdoctoral work in computer science, and now teaches and researches in computer science and engineering. The third author has a bachelors, masters, and PhD in computer science and now teaches and researches in computer science.
@@ 5. Data Flow Monitoring Results
We note that although we excluded IP addresses that were confirmed background IPs from a particular operating system or browser, in many cases, previously unseen IP addresses from the same organizations as the background traffic popped up only in the experiments on that particular OS or browser. We report these in the figures and mark them as suspected background interactions. However, we omit these suspected background contacts from the written analysis for clarity and brevity.
@@ 5.1. Idle Operating Systems
In a one hour window, Windows contacted a number of endpoints and transferred application data. Ubuntu contacted Canonical Group (the software company that maintains Ubuntu) and our university, and transferred no application data over 1000 bytes to either. Figure 2 summarizes our idle OS results.
@@ 5.2. Idle Browsers
In a one hour window, Chrome transferred application data to Google and contacted content delivery networks. Firefox transferred application data to Google Cloud and contacted content delivery networks. Brave transferred application data to Amazon. Figure 3 summarizes our idle browser results.
@@ 5.3.3. Social Media:
Instagram transferred application data to Facebook and also contacted Three Ireland Limited and Turk Telecom Times. TikTok transferred application data to Akamai: although Akamai is part of Firefox background traffic, the number of contacts to Akamai compared to the other idle experiments, as well as the volume of application data transferred to Akamai, indicate that contact to Akamai is from TikTok as well. X transferred application data to Cloudflare. Bluesky transferred application data to OVH SAS and additionally contacted a few cloud computing and content delivery networks. YouTube contacted Google and Indiana GigaPop and did not transfer any application data while running idle. Figure 6 summarizes our social media results.
@@ 5.3.4. Chat:
Slack transferred application data to Amazon and contacted content delivery networks as well as the advertising platform Rubicon Project. Discord contacted Cloudflare and Amazon, and did not transfer any application data. Figure 7 summarizes our chat application results.
@@ 5.4.1. Email:
While sending an email, Gmail contacted Google on all browsers, Amazon on Chrome, Google Cloud on Chrome, and Fastly on Brave. Of these, only Amazon on Chrome is unexpected. The University Gmail contacted Google and Amazon on all browsers. All are expected: Google is the parent company, and Amazon is expected because our university Gmail account uses single sign-on with Okta, which runs on AWS. Proton Mail contacted Proton on all browsers and Amazon on Firefox. The contact with Amazon on Firefox would be unexpected, however, Proton Mail uses AWS for alternative routing (2), thus, Proton Mail has no unexpected contact. Figure 8 summarizes our results.
@@ 5.4.2. Search:
Google Search contacted Google on all three browsers. This is expected: Google is the parent company. Duck Duck Go contacted Microsoft on all browsers. This is expected: Duck Duck Go is powered by Microsoft Bing. Brave Search contacted Amazon on all three browsers. This is expected: Brave Search runs on AWS. Figure 9 summarizes our results.
Figure 9. The organizations contacted while performing searches on Google Search, Duck Duck Go, and Brave Search. No contacts are unexpected.
@@ 5.4.3. Social Media:
Instagram contacted Facebook on all browsers; this is expected as it is the parent company. Instagram also contacted Internet service providers in other countries: Getlinks SMC Private Limited (Pakistan), Horizon Scope Mobile Telecom (Iraq), Reliance Jio Infocomm Limited (India), COSCOM Liability Limited Company (Uzbekistan), and Telecom Algeria (Algeria). This is unexpected contact given our location in the Midwestern United States. TikTok contacted Akamai, Google, and Fastly on all browsers. Akamai and Fastly are expected due to content delivery; Google is unexpected. X contacted Cloudflare, Google, and Fastly on all browsers. Cloudflare and Fastly are likely content delivery; Google is unexpected. Bluesky contacted Bluesky, OVH SAS, Datacamp Limited, and Amazon on all browsers. This is expected traffic as the parent company and content delivery networks. (Note: Bluesky used to run on Amazon Web Services; now it does not. However given that there could be some lingering contact with AWS due to platform migration, we give Bluesky the benefit of the doubt that its contact with Amazon is not due to unexpected data sharing.) Bluesky additionally contacted Amazon-AES on Brave which is likely for content delivery. YouTube contacted Google and Indiana Gigapop on all browsers. Google is the parent company of YouTube and Indiana Gigapop is a high-speed routing network for academic institutions in the Midwestern United States, so both are expected. Figure 10 summarizes our results.
@@ 5.5. Determining Levels of Control
Figure 12. Determining an operating system’s level of control based on if it sends application data during idle time, communicates with unexpected IPs, and if its parent company is a known data harvester. If any of these three criteria are met, the operating system operates at a higher level of control; if none are met, it operates at a lower level.
@@
Figure 13. Determining a browser’s level of control based on if it sends application data during idle time, communicates with unexpected IPs, and if its parent company is a known data harvester. If any of these three criteria are met, the browser operates at a higher level of control; if none are met, it operates at a lower level.
@@
Figure 14. Determining an application’s level of control based on if it sends application data during idle time, communicates with unexpected IPs, and if its parent company is a known data harvester. If any of these three criteria are met, the application operates at a higher level of control; if none are met, it operates at a lower level.
@@ 5.5.2. Browsers:
Chrome contacted Amazon which was unexpected, and transferred application data while running idle. Additionally, Chrome is owned by Google: a known data harvester (Schelter and Kunegis, 2021; Gill et al., 2013; Roesner et al., 2012; Lerner et al., 2016; Karaj et al., 2018; Englehardt and Narayanan, 2016; Krishnamurthy and Wills, 2009; Cassel et al., 2022). Thus, Chrome operates at a higher level of control.
@@ 5.5.6. Chat:
Slack transferred application data while running idle and contacted unexpected IPs, mostly advertising providers, both while running idle and while taking an action. Discord did not transfer application data while running idle; however, it did unexpectedly contact Microsoft and Google during use. Thus, both applications operate at a higher level of control.
@@ 6.1. What levels of control did participants identify as the necessary level?
Figure 15. Heat map of participant responses for if they thought user, community, or platform control of their data was necessary in different types of software for personal usage. The number inside each box indicates the number of participants who thought that level of control was necessary for that platform type. The majority of participants wanted user control for all software types. However, more participants wanted platform control for browsers, search engines, and social media.
@@
A heatmap titled ”What level of control did our participants want in personal usage?” The table has four columns: “User,” “Community,” and “Platform” under the main heading “Level of Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” and “Chat.” A legend on the bottom indicates the color shading based on the number of participants: Gray: 0 participants, Pink: 1-3 participants, Light Red: 4-7 participants, Red: 8-11 participants, Dark Red: 12-16 participants. Here is a breakdown of the data in the heatmap: OS: “User” column is dark red with a 12 indicating 12 participants responded with that. “Community” column is pink with the number 2. ”Platform” column is pink with the number 2. Browser: “User” column is red with the number 10. “Community” column is pink with the number 2. “Platform” column is light red with the number 4. Email: “User” column is dark red with the number 13. “Community” column is pink with the number 1. “Platform” column is pink with the number 2. Search: “User” column is red with the number 8. “Community” column is pink with the number 2. “Platform” column is light red with the number 6. Social Media: “User” column is red with the number 11. “Community” column is gray with the number 0. “Platform” column is light red with the number 5. Chat: “User” column is dark red with the number 16. “Community” column is gray with the number 0. “Corporate” column is gray with the number 0.
@@
Figure 16. Heat map of participant responses for if they thought user, community, or platform control of their data was necessary in different types of software for group usage. The number inside each box indicates the number of participants who thought that level of control was necessary for that platform type. The majority of participants wanted community control for all software types. However, more participants wanted platform control for search engines.
@@
A heatmap titled “What level of control did our participants want in group usage?” The table has four columns: “User,” “Community,” and “Platform” under the main heading “Level of Control,” and “Application” as the row header. The rows are “OS,” “Browser,” “Email,” “Search,” “Social Media,” and “Chat.” A legend on the bottom indicates the color shading based on the number of participants: Gray: 0 participants, Pink: 1-3 participants, Light Red: 4-7 participants, Red: 8-11 participants, Dark Red: 12-16 participants. Here is a breakdown of the data in the heatmap: OS: “User” column is gray with the number 0.“Community” column is dark red with the number 15. “Platform” column is pink with the number 1. Browser: “User” column is pink with the number 1. “Community” column is dark red with the number 12. “Platform” column is pink with the number 3. Email: “User” column is pink with the number 1. “Community” column is dark red with the number 13. “Platform” column is pink with the number 2. Search: “User” column is gray with the number 0. “Community” column is red with the number 11. “Platform” column is light red with the number 5. Social Media: “User” column is pink with the number 3. “Community” column is red with the number 10. “Platform” column is pink with the number 3. Chat: “User” column is pink with the number 2. “Community” column is dark red with the number 13. “Platform” column is pink with the number 1.